AS AD 50100 200 300 400 500 600 700 REAL OUTPUT (in bilions of dollans per year) following graph depicts a macro equilibrium. Answer the questions based on the information in the graph. (a) What is the equilibrium rate of GDP? (b) If full-employment (c) How large is the real GDP gap? real GDP is S1200, what problem does this economy have? IEshe multiplier were equal to 4, how much additional investment would be needed to increase aggregate demand by the amount of the initial GDP gap? trate the changes in autonomous investment and induced consumption that occur in ( What happens to prices (g) Is full employment restored by the AD shift? GDP gap?
Answer #1

Ans 1)

Equilibrium GDP of Economy is $500 bn dollars per year

Because equilibrium means AD equals to AS

Ans 2)

Real output Gap = actual output- possible output=500-1200=-$700 bn dollars

It is called recessionary output Gap when actual gdp is less than potential gdp.

Ans 4)

∆Y/∆I=Investment multiplier=4

We need ∆Y=700 therefore we need 700/4=175 bn dollars of investment to fill the recessionary gap